OpenStudy (anonymous):

a • b • 4 = 4 • a • b is an example of which algebraic property? Associative Property of Multiplication Transitive Property Reflexive Property Commutative Property of Multiplication

hartnn (hartnn):

A.B =B.A is commutative property. does this help ?

OpenStudy (anonymous):

Associative: (a+b) + c = a + (b+c), (ab)c = a(bc)